a golfer drives a ball toward a lake.The golf ball was struck at an angle of 27 above the horizontal with a speed of 5.6m/s.How far did the ball travel horizontally before hitting the water


1
Expert's answer
2021-10-24T18:25:45-0400

We can find the range of the ball as follows:


"R=\dfrac{v_0^2sin2\theta}{g}=\dfrac{(5.6\ \dfrac{m}{s})^2\times sin2\times27^{\circ}}{9.8\ \dfrac{m}{s^2}}=2.6\ m."